OpenStudy (anonymous):

Need a little help with my math homework. Find the Radius of a circle with the area of pi(16^2+40x+25)

OpenStudy (freckles):

Area of a circle is pir^2 \[r^2=16^2+40x+25 \]

OpenStudy (freckles):

I bet you had an x there next to 16

OpenStudy (freckles):

\[r^2=16x^2+40x+25 \] try to write the other side as something squared